d^2y/dx^2-2tanxdy/dx+3y=2secx given that y=sinx is a solution of associated homogeneous equation


1
Expert's answer
2020-07-30T15:24:24-0400

y''-2tan(x)y'+3y=2/sec(x) (1)


y=sin(x),y'=cos(x),y''=-sin(x)

y"-2tan(x)y'+3y=-sin(x)-2tan(x)cos(x)+3sin(x)=0

Really, y=sinx is a solution of associated homogeneous equation

y"-2tan(x)y'+3y=0.


Let's find the second solution y2(x) of the homogeneous equation in the form

y(x)=p(x)sin(x)

y'=p'sin(x)+pcos(x)

y''=p''sin(x)+2p'cos(x)-psin(x)

p''sin(x)+2p'cos(x)-psin(x)-2tan(x)(p'sin(x)+pcos(x))+3psin(x)=0

p''sin(x)+2p'cos(x)-psin(x)-2p'sin2(x)/cos(x)-2psin(x)+3psin(x)=0

p''sin(x)+2p'cos(x)-2p'sin2(x)/cos(x)=0

z=p'

z'sin(x)+2zcos(x)-2zsin22(x)/cos(x)=0

z'sin(x)=-2zcos(2x)/cos(x)

dz/z=-2cos(2x)/(cos(x)sin(x))dx

dz/z=-4cos(2x)/sin(2x)dx

"\int" dz/z=-4"\int"cos(2x)/sin(2x)dx

ln|z|=-2ln|sin(2x)|+C1

z=C2/sin2(2x)

p'=C2/sin2(2x)

p(x)=-C2cot(2x)/2

Let C2=-4

p(x)=2cot(2x)

y(x)=2cot(2x)sin(x)

y(x)=2cos(2x)sin(x)/sin(2x)

y(x)=2cos(2x)sin(x)/(2sin(x)cos(x))

y(x)=cos(2x)/cos(x)

So,

y2(x)=cos(2x)/cos(x)


The general solution of the homogeneous equation has the form

y(x)=C1sin(x)+C2cos(2x)/cos(x)


The equation (1) can be transform into standart form using the substitution

ln(y)=ln(z)-1/2"\int"(-2tan(x))dx

ln(y)=ln(z)-ln(cos(x))

y=z/cos(x)

z=cos(x)y

z'=-sin(x)y+cos(x)y'

z''=-cos(x)y-2sin(x)y'+cos(x)y''

cos(x)(y"-2tan(x)y'+3y)=cos(x)y''-2sin(x)y'+3cos(x)y=z''+4z

Multiplying the equation (1) by cos(x) we obtain the equation

z''+4z=2

We will seek a particular solution in the form

z=A

4A=2

A=1/2

zp(x)=1/2

yp(x)=1/(2cos(x))

The general solution of the equation (1) has the form

y(x)=C1sin(x)+C2cos(2x)/cos(x)+1/(2cos(x))